# Paint Outside the Box: Synthesizing and Selecting Training Data for Visual Grounding

### v1

@@ 4 Experiments
RefCOCO [92] 19,994 142,210 50,000
RefCOCO+ [92] 19,992 141,564 49,856
RefCOCOg [47, 50] 25,799 95,010 49,822
ReferItGame [30] 19,894 130,525 96,654
Table 2: Performance comparison with state-of-the-art baselines. All methods employ the proposed filtering scheme.
@@
GeoDiffusion [7] 47.22 34.90 28.01 18.83 25.39 24.20 29.76+1.98
Gen2Det [64] 46.61 35.19 28.85 20.65 26.10 25.00 30.40+2.62
X-Paste [96] 47.85 39.01 27.59 21.25 27.08 25.16 31.32+3.54
Ours 52.77 40.96 31.37 21.25 29.98 25.28 33.61+5.83
@@
Baselines. We compare our method with three recent data generation techniques [96, 7, 16] designed for fine-grained understanding tasks: (i) X-Paste [96] uses an ensemble of pretrained segmentation models to identify instances from generated images, which are then copied into real images. (ii) Gen2Det [64] employs a grounded editing diffusion model to edit objects in real images. (iii) GeoDiffusion [7] utilizes geometric control via text prompts to enhance text-to-image diffusion models for object detection data generation. For a fair comparison, we refrain from using the models that have been pretrained on a large scale with fine-grained annotations. Implementation details can be found in the supplementary material.
@@
Evaluation Metric. Following previous works [11, 12], we use top-1 accuracy as the evaluation metric, where a prediction is considered correct if the IoU between the predicted bounding box and the ground-truth exceeds 0.5.
@@
Implementation Details. For both the teacher and student networks, we utilize TransVG [11], one of the most popular models for visual grounding, known for being the first transformer-based architecture designed without any bells or whistles. We utilize the Stable Diffusion XL model [54] for image inpainting, with 45 denoising steps, a strength of 0.9, and a guidance scale of 7.5. BLIP ViT-L [36] is employed as the image captioner, with nucleus sampling ($p$=0.9).The hyperparameters $\lambda$ in filtering scheme is tuned using grid search on the validation set, with values of 1.0 or 0.5. For each real data point, we generate 4 synthetic images and apply our proposed selection strategy to retain one with highest score. To ensure a fair comparision, we adopt a consistent training schedule (batch size × steps = 64 × 6500) across all experiments, following previous works [95, 73, 83].
4.1 Main Results
Comparison with SOTA Baselines. Table 2 presents a comparative analysis of our proposed method against several state-of-the-art baselines across multiple benchmark datasets. We have following observations. (1) Our method outperforms the baseline trained solely on real data by an average of 5.83%, highlighting the effectiveness of our inpainting strategy and filtering scheme. (2) Models utilizing synthetic data generated by various baselines show at least a 1.98% improvement compared to the model trained exclusively on real data, demonstrating the substantial value of synthetic data in enhancing performance. (3) Our proposed method consistently outperforms all other baselines, achieving the highest accuracy across all datasets. The best-performing baseline, X-Paste, achieves an average score of 31.32%, but our approach surpasses it by 2.29%, achieving the state-of-the-art performance. Notably, our model reaches an impressive 52.77% accuracy on RefCOCO TestA and 31.37% on RefCOCO+ TestA, significantly outperforming other baselines. Additionally, our method achieves gains of 2.90% on RefCOCOg Test and 0.18% ReferIt compared to X-Paste, demonstrating robustness and generalizability across diverse datasets. We hypothesize that the relatively modest performance gain on ReferIt is due to the prevalence of small objects in this dataset. The inpainting process for such large background areas introduces noticeable artifacts, which may adversely affect model performance. In conclusion, our approach sets a new benchmark for visual grounding in our data-scarce scenario, representing a significant advancement over existing SOTA models.
@@ 4.1 Main Results
Table 3: Ablations on synthetic data and filtering scheme. $\text{Syn}_{I}$ and $\text{Syn}_{T}$ denote synthetic images and texts, respectively.
@@
Real $\text{Syn}_{I}$ $\text{Syn}_{T}$ $S_{1}$ $S_{2}$ $P$ TestA TestB TestA TestB Test Test
@@
Average Results on Three Randomly Sampled Datasets. Additionally, we sampled three different subsets from each real dataset to further evaluate performance. The results, presented in Table 2, demonstrate that despite performance variance due to the relatively small size of the sampled datasets, our method consistently outperforms the model trained solely on real data across all subsets. These findings indicate that our model’s performance is robust and generalizes well to variations in the data.
4.2 Ablation Study
@@
Ablation on Synthetic Data. We begin by evaluating the effectiveness of synthetic data by separately using generated images and texts, without any filtering scheme. We observe that using only real data results in relatively low performance, with the model achieving an average accuracy of 27.78%. Introducing synthetic images leads to a significant improvement, boosting performance by 3.53% on average, demonstrating the effectiveness of our proposed inpainting strategy. When synthetic texts are applied, the improvement is modest at 1.51%, demonstrating the importance of generated data for performance gains.
@@
Ablation on Filtering Scheme. Next, we assess the effectiveness of different components of our filtering scheme. The individual components—$S_{1}$, $S_{2}$, and $P$—result in performance gains of 1.23%, 0.71%, and 1.14%, respectively, compared to the variant without any filtering. These results demonstrate the effectiveness of each component in selecting valuable training data. Notably, the variant without the penalty term achieves an accuracy of 33.08%, outperforming all configurations that include just one filtering component. Combining synthetic data with the full filtering scheme leads to a substantial improvement, with the highest recorded gain of 1.96%. This configuration achieves the highest accuracy across all datasets, with a remarkable average accuracy of 33.61%, underscoring the effectiveness of our proposed filtering scheme.
In conclusion, all of these results demonstrate that both synthetic data and filtering schemes play crucial roles in enhancing model performance, with the combination of these elements yielding the best results.
@@
Table 4: Results obtained by isolating scores without generated text. Results with a gray background indicate performance lower than the RANDOM baseline.
Filter RefCOCO RefCOCO+ RefCOCOg
TestA TestB TestA TestB Test
RANDOM 48.52 38.84 27.47 20.27 27.34
ALL 48.45 38.68 28.40 20.29 26.90
$S_{1}$ 49.94 40.14 29.53 20.67 28.10
$S_{2}$ 49.48 39.83 27.68 20.35 29.47
$P$ 50.66 39.47 28.50 20.41 27.81
Negative Score
1-$S_{1}$ 48.17 38.90 25.98 19.25 26.40
1-$S_{2}$ 48.22 38.51 26.86 19.60 25.54
4.3 Filtering Scheme Analysis
In this section, we isolate the each component in the filtering scheme without the influence of generated text, offering insights into their impact on model performance. We create two variants of our method: (i) The RANDOM method randomly select images as training data. (ii) The ALL baseline uses all synthetic data indiscriminately for training. We have following observations from Table 4.
(1) Despite the ALL method having access to more data than the RANDOM baseline, it sometimes underperforms. For instance, in RefCOCOg, there is a performance drop of 0.44%. This observation suggests that simply increasing the amount of synthetic data does not necessarily lead to performance gains. It underscores the importance of filtering out low-quality data to ensure the effectiveness of the training data. (2) Models using any of the three filtering components outperform the two baselines. This demonstrates the critical importance of leveraging $S_{1}$ and $S_{2}$ scores to select high-quality training data, aligning with our observations in Table 3 when synthetic texts are incorporated. (3) It is worth noting that different filtering scores achieve the highest results on different datasets due to the distinct data distributions. For instance, $S_{1}$ achieves peak performance on RefCOCO+ TestA with an accuracy of 29.53%, while $S_{2}$ yields the best performance on RefCOCOg with an accuracy of 29.47%. This variability suggests that each filtering criterion captures unique aspects of the data distribution, which are beneficial for model performance in specific datasets.
Interestingly, when data is filtered using negative values of $S_{1}$ or $S_{2}$, model performance drops below the random baseline. For example, using a 1-$S_{1}$ score reduces accuracy to 48.17% on RefCOCO TestA and 25.98% on RefCOCO+ TestA, compared to the random baseline scores of 48.52% and 27.47%, respectively. Similarly, filtering with 1-$S_{2}$ results in consistent performance declines across datasets, such as a 0.61% drop on RefCOCO TestA and a 1.80% drop on RefCOCOg. The observed performance drop when using negative scores suggests that our filtering scheme is effective in identifying valuable training data, as the introduction of poorly selected data (as indicated by negative scores) negatively impacts the model’s performance.
4.4 Cross-modality Generative Models
@@
We evaluate whether POBF can seamlessly integrate with other state-of-the-art pretrained generative models. By default, POBF utilizes Stable Diffusion XL [54] as the image generator and BLIP ViT-L [36] as the image captioner. In this experiment, we test two alternative image generators: Stable Diffusion 2 [55] and Dreamshaper [^1] https://huggingface.co/Lykon/dreamshaper-8-inpainting. For image captioning, we explore replacements such as LLAVA [40] and ViT-GPT2222https://huggingface.co/nlpconnect/vit-gpt2-image-captioning. Table 5 shows the results. We observe that while the default configuration of POBF delivers robust performance, it does not always yield the highest accuracy. Specifically, model combinations using Dreamshaper as the image generator achieve superior outcomes, reaching an average accuracy of 27.63%. Furthermore, these configurations consistently outperform the baseline without our framework. These observations highlight the generalizability and effectiveness of our proposed framework.
Table 5: Performance comparison across generative models.
@@
SD 2 31.16 21.19 29.30 27.22+5.48
Dreamshaper 30.20 22.39 30.32 27.63+5.89
4.5 Ratios of Training Data
Table 6 presents the performance of our proposed framework under varying amounts of training data. We experiment with two scenarios: using 0.5% and 2.0% of the original training data. The results demonstrate that our method consistently provides substantial performance gains over the baseline. Specifically, with just 0.5% of the training data, the baseline model achieves an average accuracy of 13.94%. In contrast, our approach raises the average accuracy to 17.74%, representing a notable improvement of 3.80%. This highlights the effectiveness of our method in enhancing performance even under extremely limited data conditions. In the case of 2.0% training data, our method boosts the average accuracy by 3.99%. These consistent gains across both scenarios underscore the robustness of our approach, showcasing its potential for practical applications in data-scarce environments.
Table 6: Performance comparison across ratios of training data.
@@ 4.5 Ratios of Training Data
Real 32.97 25.11 33.77 30.62
Ours 36.54 28.84 38.47 34.61+3.99
4.6 Model Architectures
We evaluate the proposed POBF framework across different model architectures to assess its generalizability and effectiveness. Specifically, we experiment with TransVG [11], Dynamic-MDTRE [60], and QRNet [90], each of which differs in terms of the number of parameters and architectural complexity. Table 7 presents the results. Although model performance generally increases with model size, our POBF framework consistently delivers extensive performance gains across all tested architectures. Specifically, Dynamic-MDTRE exhibits an more pronounced improvement, with average accuracy rising from 23.81% to 30.84%, resulting in a gain of 7.03% points. QRNet, the most complex model with 273M parameters, also benefits from a 4.41% point increase in average accuracy. The results highlight that our method is not tailored to a specific architecture but instead offers consistent performance benefits across a diverse set of models.
Table 7: Performance comparison across model architectures. #Paras denotes the number of model parameters.

### v2

@@ 5. Experiments
RefCOCO (Yu et al., 2016) 19,994 142,210 50,000
RefCOCO+ (Yu et al., 2016) 19,992 141,564 49,856
RefCOCOg (Mao et al., 2016; Nagaraja et al., 2016) 25,799 95,010 49,822
ReferItGame (Kazemzadeh et al., 2014) 19,894 130,525 96,654
Table 2. Performance comparison with state-of-the-art baselines. All methods employ the proposed filtering scheme.
@@
GeoDiffusion (Chen et al., 2024d) 47.22 34.90 28.01 18.83 25.39 24.20 29.76+1.98
Gen2Det (Suri et al., 2024) 46.61 35.19 28.85 20.65 26.10 25.00 30.40+2.62
X-Paste (Zhao et al., 2023) 47.85 39.01 27.59 21.25 27.08 25.16 31.32+3.54
POBF (Ours) 52.77 40.96 31.37 21.25 29.98 25.28 33.61+5.83
@@
Baselines. We compare our method with three state-of-the-art baselines, all of which generate dense annotated training data. (i) X-Paste (Zhao et al., 2023) uses an ensemble of pretrained segmentation models to identify instances from generated images, which are then copied into real images. (ii) Gen2Det (Suri et al., 2024) employs a grounded editing diffusion model to edit objects in real images. (iii) GeoDiffusion (Chen et al., 2024d) utilizes geometric control via text prompts to enhance text-to-image diffusion models for direct object detection data generation. For a fair comparison, we refrain from using the models that have been pretrained on a large-scale dense annotations. Implementation details can be found in the supplementary material.
@@
Evaluation Metric. Following previous works (Deng et al., 2021, 2023), we use top-1 accuracy as the evaluation metric, where a prediction is considered correct if the IoU between the predicted bounding box and the ground-truth exceeds 0.5.
@@
Implementation Details. For both the teacher and student networks, we utilize TransVG (Deng et al., 2021), one of the most popular models for visual grounding without any bells or whistles. We utilize the Stable Diffusion XL model (Podell et al., 2023) for image inpainting, with 45 denoising steps, a strength of 0.9, and a guidance scale of 7.5. BLIP ViT-L (Li et al., 2022) is employed as the image captioner, with nucleus sampling ($p$=0.9). The hyperparameters $\lambda$ in filtering scheme is tuned using grid search on the validation set, with values of 1.0 or 0.5. For each real data point, we generate 4 synthetic images and retain one with highest score. During training, generated texts replace real captions with a probability of $q=0.3$. For a fair comparison, we adopt a consistent training schedule (batch size × steps = 64 × 6500) across all experiments, following (Zhang et al., 2023; Wang et al., 2024a; Yang et al., 2024).
5.1. Main Results
Comparison with State-of-the-Art Baselines. We have following observations from Table 2. (1) Synthetic data provides substantial performance gains. Models utilizing synthetic data generated by various baselines consistently outperform the method trained solely on real data, with at least a 1.98% improvement. This demonstrates the strong potential of synthetic data in boosting model performance. Our approach further improves upon this, achieving an average performance gain of 5.83% over the real-data-only baseline, which highlights the effectiveness of our generation strategy. (2) POBF outperforms the baseline that directly generates training data by a significant margin. GeoDiffusion is capable of directly generating images from scratch. However, it fails to correctly generate an image for batter in Figure 1. This is because fine-tuning over a fixed set of objects within a specific domain limits its generalizability to novel targets. Consequently, our method surpasses GeoDiffusion by an average margin of 3.85%. (3) POBF consistently outperforms image editing baselines. Similar to POBF, both Gen2Det and X-Paste generate training samples by editing real images. The strongest baseline, X-Paste, achieves an average score of 31.32%, while our method exceeds this by 2.29%. In Figure 1, X-Paste produces unnatural images, and Gen2Det suffers from label misalignment issues. These quality limitations hinder model performance. In contrast, POBF generates high-quality, natural images with accurately aligned bounding boxes. In addition, we hypothesize that the relatively modest improvement on ReferIt stems from the small objects in this dataset. The inpainting process requires generating very large background regions, which may introduce visible artifacts that hinder model performance. In conclusion, our method establishes a new benchmark for visual grounding in data-scarce scenarios.
@@ 5.1. Main Results
Table 3. Ablations on the synthetic data and filtering scheme. $\rm Syn_{img}$ and $\rm Syn_{txt}$ denote synthetic images and texts, respectively.
@@
Real $\rm Syn_{img}$ $\rm Syn_{txt}$ $S_{1}$ $S_{2}$ $P$ TestA TestB TestA TestB Test Test
@@
Average Results on Three Randomly Sampled Datasets. Additionally, we sampled three different subsets from each dataset to further evaluate performance in Table 2. While there is some performance variance due to the relatively small size of the sampled datasets, our method consistently outperforms the model trained solely on real data across four datasets. These results highlight the robustness of our approach to varying data distributions.
5.2. Ablation Study
@@
Ablation on Synthetic Data. We begin by evaluating the effectiveness of synthetic data by separately using generated images and texts, without the filtering scheme. We observe that using only real data results in relatively low performance, with the model achieving an average accuracy of 27.78%. Introducing synthetic images leads to a significant improvement, boosting performance by 3.53% on average, demonstrating the effectiveness of our proposed inpainting strategy. When synthetic texts are applied, the improvement is modest at 1.51%, demonstrating the importance of generated texts for performance gains.
@@
Ablation on Filtering Scheme. Next, we assess the effectiveness of different components of our filtering scheme. The individual components—$S_{1}$, $S_{2}$, and $P$—result in performance gains of 1.23%, 0.71%, and 1.14%, respectively, compared to the variant without any filtering. These results demonstrate the effectiveness of each component in selecting valuable training data. Notably, the variant excludes only the penalty term achieves an accuracy of 33.08%, surpassing all configurations that utilize a single filtering component. The full filtering scheme yields the highest accuracy across all datasets, reaching an average accuracy of 33.61%. This represents a substantial improvement of 1.96% over the baseline without filtering, further underscoring the effectiveness of our proposed filter scheme. In conclusion, both synthetic data and filtering schemes play crucial roles in enhancing model performance, with the combination of these elements yielding the best results.
5.3. Filtering Scheme Analysis
We isolate the each component in the filtering scheme, excluding the influence of generated text, offering insights into their impact on model performance. Inspired by recent advances in data selection (Xia et al., 2022; Sorscher et al., 2022; Chen et al., 2024e; Rose et al., 2023; Yang et al., 2022c; Tang et al., 2023; He et al., 2024b), we construct several adapted baselines by modifying these methods to fit within our teacher-based data selection framework. (i) Random teacher: randomly selects synthetic images as training data. (ii) No teacher: uses all synthetic data indiscriminately for training. (iii) CLIP teacher: a commonly used approach in prior works (Chen et al., 2024e; Rose et al., 2023; Yang et al., 2022c; Tang et al., 2023; He et al., 2024b), which selects synthetic images with the highest cosine similarity with the real text. (iv) Moderate-Loss (Xia et al., 2022): defines the score as the difference between a sample’s loss and the class mean loss, prioritizing samples with scores near the median. (v) Moderate-DS (Xia et al., 2022): replaces the score in Moderate-Loss with the Euclidean distance between a sample’s embedding and the class center embedding, again favoring samples near the median. (vi) Difficult-Loss (Sorscher et al., 2022): prioritizes hard examples with higher loss values. In addition, we include two variants of our method that select training samples based on the negative hardness score ($-S_{1}$) and negative overfitting score ($-S_{2}$), respectively.
We have following observations in Table 4. (1) Simply increasing the amount of synthetic data does guarantee improved performance. Although the No teacher method utilizes more data than the Random teacher baseline, it sometimes yields inferior results. For example, it shows performance drops of 0.07% and 0.16% on RefCOCO. It underscores the importance of filtering out low-quality data to ensure the effectiveness of the training data. (2) Incorporating any individual score from our filtering scheme leads to significant performance improvements over all baselines. The three methods with any individual score outperform all baselines with gains ranging from 0.58% to 1.31%. Notably, they also surpasses the CLIP teacher baseline, which is commonly used for synthetic data selection. This demonstrates the effectiveness of filtering based on hardness and overfitting perspectives compared to CLIP similarity. The observation of their effectiveness aligns with our observations in Table 3, when synthetic texts are incorporated. (3) Combining all three scores yields the best overall performance. Our full filtering scheme, which integrates all three components, outperforms all baselines with improvements ranging from 1.26% to 2.60%. In particular, it exceeds the CLIP teacher baseline by 1.32% on average. This result underscores the complementary nature of the three scores and their ability to be seamlessly integrated for enhanced performance. (4) Interestingly, filtering based on negative values of $S_{1}$ or $S_{2}$ results in worse performance than the Random teacher baseline. For example, applying $-S_{1}$ and $-S_{2}$ leads to average performance drops of 0.38% and 0.46% across the two datasets. The observed performance declines when using negative scores suggests that our filtering scheme is effective in identifying valuable training data, as the introduction of poorly selected data (as indicated by negative scores) degrades model performance.
@@ 5.3. Filtering Scheme Analysis
Table 4. Results obtained by isolating scores without generated text. Results with a gray background indicate performance lower than the Random teacher baseline.
Filter RefCOCO RefCOCO+ Avg
TestA TestB TestA TestB
Random teacher 48.52 38.84 27.47 20.27 33.76
No teacher 48.45 38.68 28.40 20.29 33.96+0.20
CLIP teacher (Chen et al., 2024e) 48.84 39.75 28.07 20.35 34.25+0.49
Moderate-Loss (Xia et al., 2022) 48.86 39.17 27.01 21.17 34.05+0.29
Moderate-DS (Xia et al., 2022) 49.61 40.47 28.21 19.84 34.31+0.55
Difficult-Loss (Sorscher et al., 2022) 46.95 37.03 27.73 19.45 32.97-0.79
Our Proposed Filtering Scheme
-$S_{1}$ 48.17 38.90 25.98 19.25 33.38-0.38
-$S_{2}$ 48.22 38.51 26.86 19.60 33.30-0.46
$S_{1}$ 49.94 40.14 29.53 20.67 35.07+1.31
$S_{2}$ 49.48 39.83 27.68 20.35 34.34+0.58
$P$ 50.66 39.47 28.50 20.41 34.67+0.91
$S_{1}$+$S_{2}$+$P$ 50.98 40.53 30.25 20.51 35.56+1.81
Relationship between 2 Scores. We visualize the relationship between the hardness score $S_{1}$ and the overfitting score $S_{2}$ on the (a) RefCOCO and (b) RefCOCOg datasets in Figure 4. As shown, there is a weak negative correlation between the two scores. The absence of a strong linear relationship between the two scores suggests that $S_{1}$ and $S_{2}$ capture different aspects of data quality, offering complementary perspectives when selecting effective training samples. To further quantify this relationship, we compute the Pearson correlation coefficients for the two datasets, which are -0.4589 for RefCOCO and -0.3810 for RefCOCOg, respectively. These moderate correlation values highlight their potential in jointly guiding data selection.
Figure 4. Scatter plot illustrating the relationship between unnormalized hardness score $S_{1}$ and overfitting score $S_{2}$ across different datasets. Each point presents a generated sample.
5.4. Cross-modality Generative Models
@@
We evaluate whether POBF can seamlessly integrate with other state-of-the-art pretrained generative models. By default, POBF utilizes Stable Diffusion XL (Podell et al., 2023) as the image generator and BLIP ViT-L (Li et al., 2022) as the image captioner. In this experiment, we test two alternative image generators: Stable Diffusion 2 (Rombach et al., 2022) and Dreamshaper [^1] https://huggingface.co/Lykon/dreamshaper-8-inpainting. For image captioning, we explore replacements such as LLAVA (Liu et al., 2024) and ViT-GPT2222https://huggingface.co/nlpconnect/vit-gpt2-image-captioning. Table 5 shows the results. We observe that while the default configuration of POBF demonstrates strong performance, it does not always yield the highest accuracy. Specifically, model combinations using Dreamshaper as the image generator achieve superior outcomes, reaching an average accuracy of 27.63%. Furthermore, these configurations consistently outperform the baseline, which highlights the generalizability and effectiveness of our proposed framework.
Table 5. Performance comparison across generative models.
@@
SD 2 (Rombach et al., 2022) 31.16 21.19 29.30 27.22+5.48
Dreamshaper1 30.20 22.39 30.32 27.63+5.89
5.5. Training Data Sizes
Table 6 presents the performance of our proposed framework under varying amounts of training data. We evaluate two scenarios: using 0.5% and 2.0% of the original training set, and report the average accuracy over three randomly sampled datasets per setting. With only 0.5% of the data, the real-data-only baseline achieves 13.97% accuracy, while our method improves this to 19.18%, yielding a substantial gain of 5.22%. At 2.0% data, our framework further increases accuracy by 5.06%. These consistent improvements highlight the robustness and effectiveness of our approach in data-scarce environments, demonstrating its potential for real-world applications where labeled data is limited.
Table 6. Performance comparison across training data sizes.
@@ 5.5. Training Data Sizes
Real 33.04 25.03 33.28 30.45
POBF (Ours) 37.68 29.88 38.97 35.51+5.06
5.6. Model Architectures
We evaluate the proposed POBF framework across different model architectures to assess its generalizability and effectiveness. Specifically, we experiment with TransVG (Deng et al., 2021), Dynamic-MDTRE (Shi et al., 2023), and QRNet (Ye et al., 2022), each of which differs in terms of the number of parameters and architectural complexity. Table 7 presents the results. Although model performance generally increases with model size, our POBF framework yields substantial improvements across all tested architectures. Notably, Dynamic-MDTRE shows the largest gain, with a 7.03% increase in average accuracy. QRNet, the most complex model with 273M parameters, also benefits from a 4.41% improvement. The results highlight that our method is not tailored to a specific architecture but instead offers consistent performance benefits across diverse model designs.
Table 7. Performance comparison across model architectures. #Paras denotes the number of trainable model parameters.